Joe has a collection of nickles and dimes that is worth $8.55. If the number of dimes is tripled and the number of nickles is decreased by 2, the value of the coins would be $23.45. how many nickles and dimes does he have. Please show your work.
Answer by scott8148(6628)

"Joe has a collection of nickles and dimes that is worth $8.55" ___ 5n + 10d = 855

" If the number of dimes is tripled and the number of nickles is decreased by 2, the value of the coins would be $23.45"
___ 5(n-2) + 10(3d) = 2345

5n - 10 + 30d = 2345

5n + 30d = 2355

subtracting 1st equation ___ 5n + 30d - (5n + 10d) = 2355 - (855)

20d = 1500

solve for d and substitute the value into any of the equations to find n